A company has 5 large buses and 6 smaller ones. A large bus holds 11 more people than the smaller one. If the capacity of all the buses is 495, how many people does each size bus hold ?
Answer by JulietG(1812)

Let x represent the capacity of the smaller bus
6x + 5(x+11) = 495
Distribute
6x + 5x + 55 = 495
11x + 55 = 495
Subtract 55 from each side
11x = 440
Divide each side by 11
x = 40
The smaller buses hold 40; therefore the larger ones hold 40+11, or 51
6 smaller buses with 40 people = 240
5 larger buses with 51 people = 255
240 + 255 = 495